OpenStudy (anonymous):

How do you write these as equasions? four less than a number three times a number the quotient of 4 and a number nine more than a number Thanks!

OpenStudy (anonymous):

x= y-4 x=3y y/4 = x x=y+9

OpenStudy (phi):

four less means -4 three times means 3 x quotient of 4 means 4/ something nine more means +9